OpenStudy (anonymous):

If a cone has a volume of 60Pi what is the radius and height?

OpenStudy (perl):

the volume of a cone is V = 1/3 h * pi * r^2

OpenStudy (anonymous):

i know that. i tried plugging the choices it gave me but i couldnt get anything im running out of time

OpenStudy (perl):

so you can substitute 60pi for volume

OpenStudy (perl):

60 pi = 1/3 * h * pi * r^2 you can divide both sides by pi and multiply both sides by 3

OpenStudy (perl):

which gives you 180 = h* r^2 an equation in two unknowns

OpenStudy (across):

The best that you can do is express the height in terms of the radius and vice versa.

OpenStudy (anonymous):

?

OpenStudy (perl):

does it say that radius = height ?

OpenStudy (across):

You can convince your instructor that you understand the problem by telling him or her that there exist infinitely-many different cones that can attain a fixed, non-trivial volume.

OpenStudy (anonymous):

nevermind. i got it. i was confusing my self

OpenStudy (anonymous):

thanks guys